The Tucson Tomato Company’s current dividend is $5. You expect the growth rate to be 0 percent for years 1 to 5, and 2 percent for years 6 to infinity. The required rate of return on this company’s equity is 10 percent. Determine the following:

A. The expected dividend at the end of year 5.

B. The expected dividend at the end of year 6.

C. The expected value of the stock at the end of year 5 (immediately after the year 5 dividend).

D. The price of the stock today
Answer #1

A. the expected dividend at the end of year 5 is :

Since there is zero growth, the expected dividend at the end of year 5 is $5.

B. At the end of year 6, the expected dividend is = $5*1.02 = $5.1

C. P5 = D6/ Re - G

= 5.1/ 0.1 - 0.02

= $63.75

D. The price of stock today is :

year 1 dividend - year 5 dividend +present value of the stock at the end of year 5

= 5/1.1^1 +5/1.1^2 + 5/1.1^3 + 5/1.1^4 + 5/1.1^5 + 63.75/1.1^5

=$4.54 + 4.13 + 3.76 + 3.42 + 3.10 + 39.58

= $58.53